OpenStudy (unklerhaukus):

Q: Express the following as existence assertions; (a) The equation x3 = 27 has a natural number solution. (b) 1,000,000 is not the largest natural number. (c) The natural number n is not a prime.

OpenStudy (unklerhaukus):

A: (a) \[\exists x(x\in\mathbb N)(x^3=27)\]

OpenStudy (unklerhaukus):

(b) \[\exists x(x\in \mathbb N)(x>10^6)\]

OpenStudy (helder_edwin):

(a) i would write \[ \large (\exists x\in\mathbb{N})(x^3=27) \] or \[ \large (\exists x)(x\in\mathbb{N}\wedge x^3=27) \]

OpenStudy (helder_edwin):

(b) \[ \large (\exists x\in\mathbb{N})(x>10^6) \] (c) \[ \large (\exists p\in\mathbb{N})(p\neq1\wedge p\neq n\wedge p\mid n) \]

OpenStudy (unklerhaukus):

(c) \[(\exists n\in \mathbb N)((\exists p,q\in\mathbb N)(pq=n) \wedge p,q\neq 1 )\]

OpenStudy (unklerhaukus):

i can't quite understand your answer for (c) @helder_edwin

OpenStudy (helder_edwin):

for n not to be prime, it has to have a divisor different from 1 and n.

OpenStudy (unklerhaukus):

does my answer for (c) work /?

OpenStudy (helder_edwin):

no.

OpenStudy (helder_edwin):

n=6 p=2 q=3

OpenStudy (unklerhaukus):

6 is not prime,

OpenStudy (helder_edwin):

but it satisfies what u wrote

OpenStudy (unklerhaukus):

isn't that what we wanted?

OpenStudy (helder_edwin):

there is a natural number for which there are two other natural numbers none of which is one and such their product is the first number.

OpenStudy (helder_edwin):

sorry. u r right. but i would write \[ \large (\exists p,q\in\mathbb{N})(pq=n\wedge p,q\neq1) \]

OpenStudy (unklerhaukus):

thank you @helder_edwin

OpenStudy (helder_edwin):

u r welcome

OpenStudy (unklerhaukus):

so these are my final answers A: (a) \[(\exists x\in\mathbb N)(x^3=27)\] (b) \[(\exists x\in \mathbb N)(x>10^6)\] (c) \[(\exists p,q\in\mathbb N)((pq=n) \wedge (p,q\neq 1 ))\]

OpenStudy (helder_edwin):

by the way. when u wrote \[ \large \exists x(x\in\mathbb{N})\color{red}{(x^3=27)} \] the red part now lies outside the range of the quantifier. there can be only one set of parenthesis after the quantifier. \[ \large (\forall x)(P(x))\color{red}{\text{whatever lies beyond this point is not linked with }\forall} \] \[ \large (\exists x)(P(x))\color{red}{\text{whatever lies beyond this point is not linked with }\exists} \]

OpenStudy (helder_edwin):

yes. those would be the final answers.

OpenStudy (unklerhaukus):

ok i think i understand

OpenStudy (helder_edwin):

great.